Optical fibers are generally composed of silica, with an index of refraction around 1.44 How fast does light travel in a silica fiber? 


1
Expert's answer
2021-08-07T03:32:41-0400
"v=\dfrac{c}{n}=\dfrac{3\cdot10^8\ \dfrac{m}{s}}{1.44}=2.08\cdot10^8\ \dfrac{m}{s}."